# Optimal Oblivious Load-Balancing for Sparse Traffic in Large-Scale Satellite Networks

### v3

@@ II-A Network Model and Notation
For any $u,v\in\{0,...,N-1\}\times\{0,...,N-1\}$, we define the “-” operation similarly. We also denote 0 to be the origin node wherever appropriate, and $v\triangleq 0-v$ for any $v\in V$. It is important to note that operators ++ and - form a closure over $V$, i.e., they form a map from $V\times V\to V$. Next, we define $e_{1}=(1,0)$ and $e_{2}=(0,1)$. Then, the set $A=\{e_{1},e_{2},-e_{1},-e_{2}\}$ denotes the set of directions along which we have links from any node $i\in V$. In other words, every node $i$ in the network has outgoing links to $\{i+e_{1},i+e_{2},i-e_{1},i-e_{2}\}$. Similarly, every node $i$ also has incoming links from $\{i+e_{1},i+e_{2},i-e_{1},i-e_{2}\}$. Therefore, we can denote every link in the network using the starting node $i$ and the direction $e\in A$ along which it is directed. More specifically, the set of links can be represented as $E=\{(i,i+e)\;|\;i\in V,e\in A\}$. Every directed edge in $E$ has capacity $c_{i,i+e}$ which we assume to be 11 for all $(i,i+e)\in E$. Finally, we use the terms “link” and “edge” interchangeably to refer to $(i,i+e)\in E$.
@@ II-D The Class of Sparse Traffic Matrices
where $\mathbf{1}({A})$ is the indicator function for condition AA. A traffic matrix $d$ is said to be $k$-sparse if it satisfies constraints (4), (5) and (6). We denote the set of all $k$-sparse traffic matrices as $\mathcal{D}^{\prime}_{k}$ defined as
@@ IV Worst-Case Load Lower Bound
The distance metric $\Delta(\cdot,\cdot)$ counts the minimum number of hops required to go from node $u$ to node $v$. We construct the Split-Diamond traffic using the distance metric above. When NN is even, the Split-Diamond traffic matrix with parameter $r$ is defined as
@@
and, $t^{*}=\left(\frac{N}{2},\frac{N}{2}\right)$ is the farthest node away from the origin. For simplicity, we only describe the case for NN even. See Figure 6 for an example. When NN is odd, the construction is similar with some minor modifications. Observe that in the above traffic matrix, every source has traffic destined only to the node farthest away from it. Intuitively, the Split-Diamond traffic matrix is a collection of locally clustered sources and destinations, where all the source-destination pairs are NN hops away. The clusters have a diamond shape, and hence the name. The parameter $r$ determines the size of this cluster. It is easy to see that the total traffic in the Split-Diamond traffic matrix is $\sum_{s,t}d^{SD(r)}_{s,s+t}=2r^{2}$. Therefore, the Split-Diamond traffic with $r\leq\sqrt{k/2}$ belongs to traffic class $\mathcal{D}_{k}$. Next, we formally state the lower bound on $\theta^{*}$ due to Split-Diamond traffic of appropriate size in Theorem 1.
@@ V Optimal Oblivious Routing for Sparse Traffic
Before we describe the LLB scheme, we define some terminology. A stem $S_{r}(i)$ of size $r$ at node $i$ is the set of nodes that differ from node $i$ only at a coordinate and are at most $r$ hops away from node ii. Mathematically, the stem is the set of nodes
@@
Lemma 4 comes from a special case of the celebrated Max-Flow Min-Cut theorem (also known as Menger’s Theorem[11]) to count the maximum number of edge-disjoint paths from a set of nodes to another set of nodes. The proof involves observing that the cut size of $S_{r}(i)$ is always $8r+4$ for $r<N/2$, and hence there must exist at least $8r$ paths from $S_{r}(0)$ to $S_{r}(t)$. Moreover, these paths can be easily found using multiple runs of breadth-first search, removing edges once a path from $S_{r}(0)$ to $S_{r}(t)$ is found successively. We defer the details to Appendix G.
@@ Appendix D Proof of Worst-Case Load Lower Bound
To prove part (a) of Theorem 1, we first consider the dual of the inner minimization problem $\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}}$ for any traffic matrix $d\in\mathcal{D}_{k}$. We focus on the case where NN is even for simplicity. The case when NN is odd is similar, with minor modifications. From strong duality, the problem $\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}$ is equivalent to
@@
where $\bar{\mathcal{P}}$ is the collection of paths from the node 0 to every $t\in V\backslash\{0\}$. From the above formulation, observe that $\theta^{*}$ involves computing the maximum possible minimum sum cost of routing from 0 to every node $t$. Define $B_{r}(i)\triangleq\{j\in V|\Delta(i,j)\leq r-1\}$ to be the ball around node $i$ containing nodes which are at most r1r-1 hops away from ii. Observe that for a sufficiently “distant” node $t$, any route from 0 to $t$ must traverse through $B_{r}(0)$ and $B_{r}(t)$ for a sufficiently small $r$. In fact, it must must traverse through at least $r$ edges emanating from nodes in $B_{r}(0)$ that are directed away from the origin, and traverse through at least $r$ edges emanating from nodes in $B_{r}(t)$ that are directed toward $t$. This idea is visualized in Figure 9.
@@
The split diamond traffic matrix of size $r$ does exactly this. It selects $t$ to be the farthest node from the origin, which is $(N/2,N/2)$, only when $k\leq N^{2}/2$. The sources and destinations of traffic are carefully selected such that every hop that is directed away from the origin in $B_{r}(0)$ incurs a cost of 11 and every hop that is directed toward $t$ in $B_{r}(t)$ incurs a cost of 11. Hence, $\lambda=1$ and every path from 0 to $t=(N/2,N/2)$ incurs a cost of $2r$. As mentioned earlier, for the split diamond traffic of size $r$, $\sum_{s,t}d^{SD(r)}_{s,s+t}=2r^{2}$ must be lesser than or equal to $k$ for $d^{SD(r)}$ to belong to $\mathcal{D}_{k}$. The sum cost of the objective under Split-Diamond traffic of size $r$ is $\frac{1}{4}\times 2r=\frac{r}{2}$. We maximize this sum cost under the constraint $2r^{2}\leq k,r\in\mathbb{N}$. When $2k$ is a perfect square, the best choice of $r$ would be $\sqrt{k/2}$ and the value of the objective is $\sqrt{2k}/4$. Since $d^{SD(r)}$ is a feasible solution, we arrive at $\theta^{*}\geq\sqrt{2k}/4$ when $2k$ is a perfect square.
@@ Appendix E Optimality of Local Load-Balancing Policy
Appendix F Modifications to LLB policy under specific conditions
As discussed in Section V, it suffices to describe the routing policy when $t_{x}\leq t_{y}\leq N/2$ as for other destinations, the routing can be obtained from automorphism transformations. In this section, we limit ourselves to $t_{x}\leq t_{y}\leq N/2$ for ease of presentation. (For modifications to the GLLB algorithm, just replace $S_{r}(0)$ and $S_{r}(t)$ with $S_{r_{1},r_{2}}(0)$ and $S_{r_{1},r_{2}}(t)$ respectively. The following arguments are valid for the GLLB routing algorithm as well).
@@ Appendix F Modifications to LLB policy under specific conditions
When there is overlap between the stems $S_{r}(0)$ and $S_{r}(t)$, the above described routing policy cannot be used as is. If there are nodes $j$ such that $j\in S_{r}(0)\cap S_{r}(t)$, there needs to be minor modifications in the routing scheme. This happens when $t_{x}\leq r$ or $t_{y}\leq r$. Without loss in generality, there can only be the two types of overlap between stems, Type-I and type-II, as shown in Fig. 10 and Fig. 11.
F-1 Type-I Overlap
In Type-I overlap, the stems $S_{r}(0)$ and $S_{r}(t)$ overlap only along one of the four legs of the stem (Fig. 10). Under Type-I overlap, we define modified stems $S^{I}_{r}(0)$ and $S^{I}_{r}(t)$,
$S^{I}_{r}(0)$ $=\{j\in V|i\in S_{r}(0)\text{ and }\Delta(0,j)\leq\Delta(t,j)\}$
$S^{I}_{r}(t)$ $=\{j\in V|i\in S_{r}(t)\text{ and }\Delta(t,j)\leq\Delta(0,j)\}.$
F-2 Type-II Overlap
In Type-II overlap, the stems $S_{r}(0)$ and $S_{r}(t)$ overlap exactly at two nodes at two of the four legs of the stem respectively (Fig. 11). Under Type-II overlap, let $(0,h_{x})$ and $(h_{y},0)$ be the indices of the nodes that lie in both $S_{r}(0)$ and $S_{r}(t)$. We define modified stems $S^{II}_{r}(0)$ and $S^{II}_{r}(t)$,
$S^{II}_{r}(0)$ $=\{j\in V|\;j\in S_{r}(0)\text{ and }j_{x}\leq h_{x}\text{ and }j_{y}\leq h_{y}\}$
$S^{II}_{r}(t)$ $=\{j\in V|\;j\in S_{r}(t)\text{ and }j_{x}\leq t_{x}-h_{x}\text{ and }j_{y}\leq t_{y}-h_{y}\}.$
@@ F-2 Type-II Overlap
(a) First phase - Distribute source traffic among stem nodes along the shortest path. The source (origin) is highlighted in dark red and the stem nodes $S^{I}_{r}(0)$ are shaded with red vertical lines. The boxed nodes were common to both the initial stems $S_{r}(0)$ and $S_{r}(t)$. The checkered node belongs to both $S^{I}_{r}(0)$ and $S^{I}_{r}(t)$.
@@
(b) Second phase - Route traffic from $S^{I}_{r}(0)$ to $S^{I}_{r}(t)$ along edge-disjoint paths with two edge-disjoint paths emanating from each node in $S^{I}_{r}(0)$ and terminating at each node in $S^{I}_{t}(t)$. The last node in the stem $S^{I}_{r}(0)$ along the overlapping leg also carries residual traffic to the last node in stem $S^{I}_{r}(t)$, which is indicated by an asterisk.
@@
(c) Third phase - Aggregate traffic from the stem nodes $S^{I}_{r}(t)$ to the destination node $t$ along the shortest path. The destination node is highlighted in dark blue and the stem nodes $S^{I}_{r}(t)$ are shaded with blue horizontal lines. The boxed nodes were common to both the initial stems $S_{r}(0)$ and $S_{r}(t)$. The checkered node belongs to both $S^{I}_{r}(0)$ and $S^{I}_{r}(t)$.
Figure 10: LLB Routing Scheme with parameter $r=2$ with destination $t=(0,3)$ under Type-I overlap in a $7\times 7$ torus.
@@
(a) First phase - Distribute source traffic among stem nodes along the shortest path. The source (origin) is highlighted in dark red and the stem nodes $S^{II}_{r}(0)$ are shaded with red vertical lines. The boxed nodes were common to both the initial stems $S_{r}(0)$ and $S_{r}(t)$. The checkered node belongs to both $S^{II}_{r}(0)$ and $S^{II}_{r}(t)$.
@@
(b) Second phase - Route traffic from $S^{II}_{r}(0)$ to $S^{II}_{r}(t)$ along edge-disjoint paths with two edge-disjoint paths emanating from each node in $S^{II}_{r}(0)$ and terminating at each node in $S^{II}_{t}(t)$. There is no residual traffic that needs to be carried in this case
@@
(c) Third phase - Aggregate traffic from the stem nodes $S^{II}_{r}(t)$ to the destination node $t$ along the shortest path. The destination node is highlighted in dark blue and the stem nodes $S^{II}_{r}(t)$ are shaded with blue horizontal lines. The boxed nodes were common to both the initial stems $S_{r}(0)$ and $S_{r}(t)$. The checkered node belongs to both $S^{II}_{r}(0)$ and $S^{II}_{r}(t)$.
Figure 11: LLB Routing Scheme with parameter $r=2$ with destination $t=(2,1)$ under Type-II overlap in a $7\times 7$ torus.
Once we define the modified stem for each type of overlap, the modifications in each routing phase is fairly straightforward. Henceforth, $S^{\prime}_{r}(0)$ and $S^{\prime}_{r}(t)$ refer to the modified stems in each case of overlap.
$\bullet$ Modification in the first phase: Distribute $1/4r$ fraction of traffic to each of the nodes in the stems $S^{\prime}_{r}(0)$, except the last node in each of the legs. If the node is the last node in its leg and is $h$ hops away, it receives (rh+1)/4r(r-h+1)/4r fraction of the traffic.
@@
$\bullet$ Modification in the second phase: Route traffic from each node in the stem $S^{\prime}_{r}(0)$ to each node in the stem $S^{\prime}_{r}(t)$ (or $S^{\prime}_{r}(t)$ along edge-disjoint paths, such that each node in $S^{\prime}_{r}(0)$ sends equal traffic along 2 unique edge-disjoint paths and each node in $S^{\prime}_{r}(t)$ receives traffic along 2 edge-disjoint paths. Each path carries $1/8r$ fracton of traffic. If there is residual traffic at a node in $S^{\prime}_{r}(0)$, then there must exist a node neighboring it that is in $S^{\prime}_{r}(t)$. The residual traffic must be directly sent to that neighbor. If a node belongs to both $S^{\prime}_{r}(0)$ and $S^{\prime}_{r}(t)$, then the node needs to do nothing in the second phase.
@@
$\bullet$ Modification in the third phase: Aggregate traffic from the nodes in stem $S^{\prime}_{r}(t)$ to the final destination of the traffic, $t$, along the shortest path. As a consequence of the first two stages, $1/4r$ fraction of traffic needs to be aggregated from each node in $S^{\prime}_{r}(t)$ except the last nodes in each leg. A fraction of (rh+1)/4r(r-h+1)/4r of the traffic will be aggregated from the last node in each leg, where the last node is $h$ hops away.

### v4

@@ II-A Network Model and Notation
For any $u,v\in\{0,...,N-1\}\times\{0,...,N-1\}$, we define the “-” operation similarly. We also denote 0 to be the origin node wherever appropriate, and $v\triangleq 0-v$ for any $v\in V$. It is important to note that operators ++ and - form a closure over $V$, i.e., they form a map from $V\times V\to V$. Next, we define $e_{1}=(1,0)$ and $e_{2}=(0,1)$. Then, the set $A=\{e_{1},e_{2},-e_{1},-e_{2}\}$ denotes the set of directions along which we have links from any node $i\in V$. In other words, every node $i$ in the network has outgoing links to $\{i+e_{1},i+e_{2},i-e_{1},i-e_{2}\}$. Similarly, every node $i$ also has incoming links from $\{i+e_{1},i+e_{2},i-e_{1},i-e_{2}\}$. Therefore, we can denote every link in the network using the starting node $i$ and the direction $e\in A$ along which it is directed. More specifically, the set of links can be represented as $E=\{(i,i+e)\;|\;i\in V,e\in A\}$. Every directed edge in $E$ has capacity $c_{i,i+e}$ which we assume to be $1$ for all $(i,i+e)\in E$. Finally, we use the terms “link” and “edge” interchangeably to refer to $(i,i+e)\in E$.
@@ II-D The Class of Sparse Traffic Matrices
where $\mathbf{1}({A})$ is the indicator function for condition $A$. A traffic matrix $d$ is said to be $k$-sparse if it satisfies constraints (4), (5) and (6). We denote the set of all $k$-sparse traffic matrices as $\mathcal{D}^{\prime}_{k}$ defined as
@@ IV Worst-Case Load Lower Bound
The distance metric $\Delta(\cdot,\cdot)$ counts the minimum number of hops required to go from node $u$ to node $v$. We construct the Split-Diamond traffic using the distance metric above. When $N$ is even, the Split-Diamond traffic matrix with parameter $r$ is defined as
@@
and, $t^{*}=\left(\frac{N}{2},\frac{N}{2}\right)$ is the farthest node away from the origin. For simplicity, we only describe the case for $N$ even. See Figure 6 for an example. When $N$ is odd, the construction is similar with some minor modifications. Observe that in the above traffic matrix, every source has traffic destined only to the node farthest away from it. Intuitively, the Split-Diamond traffic matrix is a collection of locally clustered sources and destinations, where all the source-destination pairs are $N$ hops away. The clusters have a diamond shape, and hence the name. The parameter $r$ determines the size of this cluster. It is easy to see that the total traffic in the Split-Diamond traffic matrix is $\sum_{s,t}d^{SD(r)}_{s,s+t}=2r^{2}$. Therefore, the Split-Diamond traffic with $r\leq\sqrt{k/2}$ belongs to traffic class $\mathcal{D}_{k}$. Next, we formally state the lower bound on $\theta^{*}$ due to Split-Diamond traffic of appropriate size in Theorem 1.
@@ V Optimal Oblivious Routing for Sparse Traffic
Before we describe the LLB scheme, we define some terminology. A stem $S_{r}(i)$ of size $r$ at node $i$ is the set of nodes that differ from node $i$ only at a coordinate and are at most $r$ hops away from node $i$. Mathematically, the stem is the set of nodes
@@
Lemma 4 comes from a special case of the celebrated Max-Flow Min-Cut theorem (also known as Menger’s Theorem[11]) to count the maximum number of edge-disjoint paths from a set of nodes to another set of nodes. The proof involves observing that the cut size of $S_{r}(i)$ is always $8r+4$ for $r<N/2$, and hence there must exist at least $8r$ paths from $S_{r}(0)$ to $S_{r}(t)$. Moreover, these paths can be easily found using multiple runs of breadth-first search (or a Max-Flow algorithm), removing edges once a path from $S_{r}(0)$ to $S_{r}(t)$ is found successively. We defer the details to Appendix G.
@@ Appendix D Proof of Worst-Case Load Lower Bound
To prove part (a) of Theorem 1, we first consider the dual of the inner minimization problem $\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}}$ for any traffic matrix $d\in\mathcal{D}_{k}$. We focus on the case where $N$ is even for simplicity. The case when $N$ is odd is similar, with minor modifications. From strong duality, the problem $\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}$ is equivalent to
@@
where $\bar{\mathcal{P}}$ is the collection of paths from the node 0 to every $t\in V\backslash\{0\}$. From the above formulation, observe that $\theta^{*}$ involves computing the maximum possible minimum sum cost of routing from 0 to every node $t$. Define $B_{r}(i)\triangleq\{j\in V|\Delta(i,j)\leq r-1\}$ to be the ball around node $i$ containing nodes which are at most r1r-1 hops away from $i$. Observe that for a sufficiently “distant” node $t$, any route from 0 to $t$ must traverse through $B_{r}(0)$ and $B_{r}(t)$ for a sufficiently small $r$. In fact, it must must traverse through at least $r$ edges emanating from nodes in $B_{r}(0)$ that are directed away from the origin, and traverse through at least $r$ edges emanating from nodes in $B_{r}(t)$ that are directed toward $t$. This idea is visualized in Figure 9.
@@
The split diamond traffic matrix of size $r$ does exactly this. It selects $t$ to be the farthest node from the origin, which is $(N/2,N/2)$, only when $k\leq N^{2}/2$. The sources and destinations of traffic are carefully selected such that every hop that is directed away from the origin in $B_{r}(0)$ incurs a cost of $1$ and every hop that is directed toward $t$ in $B_{r}(t)$ incurs a cost of $1$. Hence, $\lambda=1$ and every path from 0 to $t=(N/2,N/2)$ incurs a cost of $2r$. As mentioned earlier, for the split diamond traffic of size $r$, $\sum_{s,t}d^{SD(r)}_{s,s+t}=2r^{2}$ must be lesser than or equal to $k$ for $d^{SD(r)}$ to belong to $\mathcal{D}_{k}$. The sum cost of the objective under Split-Diamond traffic of size $r$ is $\frac{1}{4}\times 2r=\frac{r}{2}$. We maximize this sum cost under the constraint $2r^{2}\leq k,r\in\mathbb{N}$. When $2k$ is a perfect square, the best choice of $r$ would be $\sqrt{k/2}$ and the value of the objective is $\sqrt{2k}/4$. Since $d^{SD(r)}$ is a feasible solution, we arrive at $\theta^{*}\geq\sqrt{2k}/4$ when $2k$ is a perfect square.
@@ Appendix E Optimality of Local Load-Balancing Policy
Appendix F LLB When Source and Destination Stems Overlap
@@
When there is overlap between the stems $S_{r}(0)$ and $S_{r}(t)$, the above described routing policy cannot be used as is. If there are nodes $j$ such that $j\in S_{r}(0)\cap S_{r}(t)$, there needs to be minor modifications in the routing scheme. This happens when $t_{x}\leq r$ or $t_{y}\leq r$. Observe that there are nodes that are both in the source stem $S_{r}(0)$ and the destination stem $S_{r}(t)$.
@@
(a) First phase - Distribute source traffic among stem nodes along the shortest path. The source (origin) is highlighted in dark red and the stem nodes $S_{r}(0)$ are shaded with red vertical lines. The boxed nodes are common to both the stems $S_{r}(0)$ and $S_{r}(t)$.
@@
(b) Second phase - Route traffic from $S_{r}(0)$ to $S_{r}(t)$ along edge-disjoint paths with two edge-disjoint paths emanating from each node in $S_{r}(0)$ and terminating at each node in $S_{t}(t)$. If a node in $S_{r}(0)$ also belongs to $S_{r}(t)$, then we do nothing.
@@
(c) Third phase - Aggregate traffic from the stem nodes $S_{r}(t)$ to the destination node $t$ along the shortest path. The destination node is highlighted in dark blue and the stem nodes $S_{r}(t)$ are shaded with blue horizontal lines. The boxed nodes were common to both the initial stems $S_{r}(0)$ and $S_{r}(t)$.
Figure 10: LLB Routing Scheme with parameter $r=2$ with destination $t=(0,3)$ in a $7\times 7$ torus. Observe that the final flow from source to destination can have loops or redundant edges. These loops or redundant edges can be removed by revising the net flow in the forward direction.
@@
(a) First phase - Distribute source traffic among stem nodes along the shortest path. The source (origin) is highlighted in dark red and the stem nodes $S_{r}(0)$ are shaded with red vertical lines. The boxed nodes are common to both the stems $S_{r}(0)$ and $S_{r}(t)$.
@@
(b) Second phase - Route traffic from $S_{r}(0)$ to $S_{r}(t)$ along edge-disjoint paths with two edge-disjoint paths emanating from each node in $S_{r}(0)$ and terminating at each node in $S_{t}(t)$. If a node in $S_{r}(0)$ also belongs to $S_{r}(t)$, then we do nothing.
@@
(c) Third phase - Aggregate traffic from the stem nodes $S_{r}(t)$ to the destination node $t$ along the shortest path. The destination node is highlighted in dark blue and the stem nodes $S_{r}(t)$ are shaded with blue horizontal lines. The boxed nodes were common to both the initial stems $S_{r}(0)$ and $S_{r}(t)$.
Figure 11: LLB Routing Scheme with parameter $r=2$ with destination $t=(2,1)$ in a $7\times 7$ torus. Observe that the final flow from source to destination can have loops or redundant edges. These loops or redundant edges can be removed by revising the net flow in the forward direction.
We keep track of these nodes and run the three phases of the LLB policy as earlier. The phases are described below.
$\bullet$ First phase: Distribute $1/4r$ fraction of traffic to each of the nodes in the stems $S_{r}(0)$ along the shortest path. The shortest path is clearly along the legs of the stem, along edges that move away from the source 0.
@@
$\bullet$ Second phase: Route traffic from each node in the stem $S_{r}(0)$ to each node in the stem $S_{r}(t)$ along edge-disjoint paths, such that each node in $S_{r}(0)$ sends equal traffic along 2 unique edge-disjoint paths and each node in $S_{r}(t)$ receives traffic along 2 edge-disjoint paths. If a node is in both $S_{r}(0)$ and $S_{r}(t)$, no traffic needs to be explicitly forwarded from that node, as it is already in the destination stem. Additionally, the paths in the second phase must not use the edges used in the first and third phases. Each path carries $1/8r$ fraction of traffic.
@@
$\bullet$ Third phase: Aggregate traffic from the nodes in stem $S_{r}(t)$ to the final destination of the traffic, $t$, along shortest paths. As a consequence of the first two stages, $1/4r$ fraction of traffic needs to be aggregated from each node in $S_{r}(t)$. The shortest path is clearly along the legs of the stem, along edges that move toward the destination $t$.
Due to overlaps in the stem sets, there could be loopy paths that could formed. There could also be edges used in the first and third phase that are used twice (once in the first phase and once in the third phase). As a final step, the final flow in the forward direction needs to be updated to remove loops or redundant use of edges. Figures 10 and 11 display examples of the LLB policy when the stems overlap.